# The equilibrium properties of obvious strategy profiles

### v1

@@ 4.3 Related literature
Various progress has been made in large game theory since the foundational papers Schmeidler (1973) and Mas-Colell (1984). Recent developments in large game theory include Kalai (2004), Sun (2006), Khan et al. (2013), Qiao and Yu (2014), Yu (2014), He et al. (2017), Khan et al. (2017), Kalai and Shmaya (2018), Carmona and Podczeck (2020), Khan et al. (2020), Sun et al. (2020), Hellwig (2022), Anderson et al. (2022a, b), and Yang (2022). Beyond theoretical advancements, the large game model is also widely used and studied in applied theory literature. Such papers include Acemoglu and Jensen (2015) in macroeconomics, Yannelis (2009) in large economy, Olszewski and Siegel (2016) in large contest games, Che et al. (2019) in large matching, Barelli et al. (2022) in large elections, and Parise and Ozdaglar (2023) in large network games. We will review the directly related papers in this subsection.
The research most closely related to this paper is by Kalai and Shmaya (2018), who studied a model of incomplete-information repeated games with a large but finite number of players. Given the model’s complexity, they introduced the concept of a (symmetric) imagined-continuum equilibrium to simplify the analysis. In this framework, the societal summary—referred to as the “macro-strategy”—is defined as the arithmetic average of players’ strategies, and individual deviations do not impact this societal summary. Their key findings demonstrated the existence of a symmetric Markov equilibrium in the imagined-continuum game and showed that this equilibrium leads to an approximate equilibrium in the finite-player setting, given that a Lipschitz continuity condition on outcome generating function is satisfied. The key distinctions between Kalai and Shmaya (2018) and this paper are as follows.
First, while both Kalai and Shmaya (2018) and this paper focus on symmetric equilibria in the large continuum game, their motivations differ. Kalai and Shmaya (2018) used symmetry primarily to simplify the characterization of equilibrium, whereas we adopt symmetry to address coordination issues arising from the implementation of the large equilibrium outcome.
Second, in Kalai and Shmaya (2018), players possess the same characteristic (i.e., identical payoff function), resulting in the same strategy being adopted in a symmetric equilibrium in both the continuum and finite games. In this context, the continuity condition required in Theorem 1 is trivially satisfied. In contrast, our model allows for different players to have distinct characteristics. As illustrated in Example 1, the continuity condition on the auxiliary function becomes indispensable in our framework.
Third, the proof of the main results in Kalai and Shmaya (2018) technically relies on a finite action set, whereas our proof accommodates a general action space. Moreover, the payoff function in our model is not necessarily Lipschitz continuous.
@@
This paper builds on prior research concerning symmetric equilibrium in large games, including works by Mas-Colell (1984), Khan and Sun (1995a, b), and Sun et al. (2020). Mas-Colell (1984) established the existence of pure strategy symmetric equilibria in large games with finite action sets and atomless characteristic distributions (see Definition 8). Khan and Sun (1995a, b) expanded on this result, addressing large games with countable action sets. More recently, Sun et al. (2020) explored randomized strategy symmetric equilibria and showed the existence of symmetric equilibria in general large games. These foundational existence results in the literature underpin our main contribution. A key distinction in our study is its focus on a finite number of players, whereas existing research predominantly examines large games involving a continuum of players.
@@ 5.1 Technical preparations
Let $A$ be a compact metric space endowed with its Borel $\sigma$-algebra $\mathcal{B}(A)$, and $d_{A}$ denote the metric on $A$. In this subsection, we introduce two equivalent metrics on $\mathcal{M}(A)$. Based on such metrics, we show that the distance between the sequence of realized society summaries and the sequence of society summaries converges to 0 in probability.
@@
It is known in the literature that $\rho$ and $\beta$ are equivalent metrics; see, for example, Bogachev (2007, Theorem 8.3.2). Given a sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ where each $g^{n}$ is a randomized strategy profile of game $G^{n}$, let $\{x^{n}_{i}\}_{i\in I^{n},n\in\mathbb{Z}_{+}}$ be a sequence of random variables mapping from a probability space $(\Omega,\Sigma,\mathbb{P})$ to $A$ such that
@@
    for each $i\in I^{n}$ and $n\in\mathbb{Z}_{+}$, the distribution induced from $x^{n}_{i}$ is $g^{n}(i)$;
@@
    for each $n\in\mathbb{Z}_{+}$, the random variables $\{x^{n}_{i}\}_{i\in I^{n}}$ are pairwise independent.
@@
Let $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite-player games, and $g^{n}$ be a randomized strategy profile of $G^{n}$ for each $n\in\mathbb{Z}_{+}$. For each $\omega\in\Omega$, let $s(x^{n})(\omega)={\sum_{i\in I^{n}}{{\delta_{{x^{n}_{i}(\omega)}}\lambda^{n}(i)}}}$ be a realized societal summary of the randomized strategy profile $g^{n}$, hence $s(x^{n})$ can be viewed as a random variable from $(\Omega,\Sigma,\mathbb{P})$ to $\mathcal{M}(A)$. Then we have
$\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0\text{ and }\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0\text{ in probability,}$,
where $s(g^{n})=\int_{I^{n}}g^{n}(i)\mathrm{d}\lambda^{n}(i)$.
@@
We divide the proof into two steps. In step 1, we show that for any bounded continuous function $h\colon A\to\mathbb{R}$ with $||h||_{BL}\leq 1$, $\int_{A}h{\rm{d}}\bigl{(}{s(x^{n})-s(g^{n})}\bigr{)}\to 0$ in probability. In step 2, we show that $\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability. Finally, by the equivalence of $\rho$ and $\beta$, we obtain that $\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability.
@@
$\int_{A}h{\rm{d}}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}=\sum\limits_{i\in I^{n}}\lambda^{n}(i)\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}-\mathbb{E}\bigl{[}h\bigl{(}x^{n}_{i}\bigr{)}\bigr{]}\Bigr{)}\to 0$ (1)
in probability. Fix any $n\in\mathbb{Z}_{+}$, since $\{x^{n}_{i}\}_{i\in I^{n}}$ are pairwise independent and $h$ is a bounded continuous function, we know that $\{h(x^{n}_{i})\}_{i\in I^{n}}$ are also pairwise independent. By the definition of $\|h\|_{BL}\leq 1$, we have $\|h\|_{\infty}\leq 1$ and hence $-1\leq h(x^{n}_{i})\leq 1$, $\text{var}\bigl{(}h(x^{n}_{i})\bigr{)}\leq 1$, for all $i\in I^{n}$. Moreover, by the independence of $\{h(x^{n}_{i})\}_{i\in I^{n}}$, we have
$\text{var}\Bigl{(}\sum\limits_{i\in I^{n}}h(x^{n}_{i})\lambda^{n}(i)\Bigr{)}=\sum\limits_{i\in I^{n}}\bigl{(}\lambda^{n}(i)\bigr{)}^{2}\text{var}\bigl{(}h(x^{n}_{i})\bigr{)}.$
Since $\mathbb{E}\bigl{[}\sum\limits_{i\in I^{n}}\lambda^{n}(i)h(x^{n}_{i})\bigr{]}=\sum\limits_{i\in I^{n}}\lambda^{n}(i)\mathbb{E}\bigl{[}h(x^{n}_{i})\bigr{]}$, for any $\varepsilon>0$, we have
$\mathbb{P}\Bigl{(}\Bigl{|}\sum_{i\in I^{n}}\lambda^{n}(i)\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}-\mathbb{E}\bigl{[}h\bigl{(}x^{n}_{i}\bigr{)}\big{]}\Bigr{)}\Bigr{|}\leq\varepsilon\Bigr{)}$
$\geqslant 1-\frac{\sum\limits_{i\in I^{n}}\bigl{(}\lambda^{n}(i)\bigr{)}^{2}\text{var}\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}\Bigr{)}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\sum\limits_{i\in I^{n}}\bigl{(}\lambda^{n}(i)\bigr{)}^{2}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\mathop{\sup}\limits_{j\in I^{n}}\lambda^{n}(j)\sum\limits_{i\in I^{n}}{\lambda^{n}(i)}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\mathop{\sup}\limits_{j\in I^{n}}\lambda^{n}(j)}{\varepsilon^{2}},$ (2)
where the first inequality is due to the Chebyshev’s inequality, and the last inequality follows from the fact that $\sum_{i\in I^{n}}{\lambda^{n}(i)}=1$. Combining with $\sup_{j\in I^{n}}\lambda^{n}(j)\to 0$ as $n\to\infty$, we can finish the proof of Formula (1).
@@
$\sum\limits_{i\in I^{n}}\lambda^{n}(i)\Bigl{(}p_{l}(x^{n}_{i})-\mathbb{E}\bigl{[}p_{l}(x^{n}_{i})\bigr{]}\Bigr{)}\rightarrow 0$ (3)
@@ 5.2 Proof of Lemma 1
Let $s(G,g)=\int_{I}\delta_{G(i)}\otimes g(i)\operatorname{d\!}\lambda(i)$ be the joint distribution of $G$ and $g$ on the product space $\mathcal{U}_{A}\times A$. Clearly, the marginal distribution $s(G,g)|_{\mathcal{U}_{A}}=\lambda G^{-1}$. Since both $\mathcal{U}_{A}$ and $A$ are Polish spaces, there exists a family of Borel probability measures $\{\mathcal{S}(u,\cdot)\}_{u\in\mathcal{U}_{A}}$ in $\mathcal{M}(A)$, which represents the disintegration of $s(G,g)$ with respect to $\lambda G^{-1}$ on $\mathcal{U}_{A}$. Let $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ be defined such that $\overline{g}(u)=\mathcal{S}(u)$ for all $u\in\mathcal{U}_{A}$. According to Sun et al. (2020, Lemma 5), the composition mapping $\widetilde{g}=\overline{g}\circ G$ is a Nash equilibrium of $G$ that satisfies $s(\widetilde{g})=s(g)=\tau$. Furthermore, we can ensure that for any player $i\in I$, her strategy $\widetilde{g}(i)$ is a best response with respect to the society summary $s(\widetilde{g})$. This requirement can be met by modifying the strategies of a subset of players with measure zero.
@@ 5.3 Proof of Theorem 1
The major difficulty of this proof is to estimate the difference between $u_{i}^{n}(g^{n})$ and $u_{i}^{n}(\mu,g_{-i}^{n})$ for all $\mu\in\mathcal{M}(A)$. We divide this estimation into five steps. Step 1 shows that we can focus on a subset of payoff functions that consists of equicontinuous and uniformly bounded functions. We then estimate the difference between $u_{i}^{n}(\mu,g^{n}_{-i})$ and $u^{n}_{i}\bigl{(}\mu,s(\mu,g^{n}_{-i})\bigr{)}$ for all $\mu\in\mathcal{M}(A)$ in step 2. Such an estimation result enables us to estimate the difference between $u_{i}^{n}(g^{n})$ and $u_{i}^{n}\bigl{(}g^{n}_{i},s(\widetilde{g})\bigr{)}$ in step 3, and the difference between $u_{i}^{n}(\mu,g_{-i}^{n})$ and $u_{i}^{n}\bigl{(}\mu,s(\widetilde{g})\bigr{)}$ for all $\mu\in\mathcal{M}(A)$ in step 4. Finally in step 5, we combine step 4 and step 5 to finish the proof.
Step 1. We show that for any $\varepsilon>0$, there exist a sequence of subsets $S^{n}\subseteq I^{n}$ such that $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$, and $\{u_{i}^{n}\}_{i\in S^{n},n\in\mathbb{Z}_{+}}$ are equicontinuous and uniformly bounded by a constant $M_{\varepsilon}$. For simplicity, let $\mathcal{W}^{n}=\lambda^{n}(G^{n})^{-1}$, and $\mathcal{W}=\lambda G^{-1}$. Since $A\times\mathcal{M}(A)$ is a compact metric space, the space of bounded and continuous functions $\mathcal{U}_{A}$ on $A\times\mathcal{M}(A)$ is a Polish space. By using the Prohorov theorem (Billingsley (1999, Theorem 5.2)), we know that $\{\mathcal{W}^{n}\}_{n\in\mathbb{Z}_{+}}$ is tight, which means that for any $\varepsilon>0$, there exists a compact set $K_{\varepsilon}\subset\mathcal{U}_{A}$ such that $\mathcal{W}^{n}(K_{\varepsilon})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$.
Since $K_{\varepsilon}$ is a compact set that consists of bounded and continuous functions, the Arzelà-Ascoli theorem (Munkres (2000, Theorem 45.4)) implies that all the functions in $K_{\varepsilon}$ are equicontinuous and uniformly bounded. Let $M_{\varepsilon}$ denote a bound of all the functions in $K_{\varepsilon}$, and $S^{n}=\{i\in I^{n}|u_{i}^{n}\in K_{\varepsilon}\}$ for all $n\in\mathbb{Z}_{+}$. It is clear that $\lambda^{n}(S^{n})>1-\frac{\varepsilon}{2}$.
Step 2. We estimate $\Bigl{|}u_{i}^{n}(\mu,g^{n}_{-i})-u^{n}_{i}\bigl{(}\mu,s(\mu,g^{n}_{-i})\bigr{)}\Big{|}$ in this step. To be precise, we prove that for any $\varepsilon>0$ and any sequence of randomized strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$, there exists $N_{\varepsilon}\in\mathbb{Z}_{+}$ such that for all $n\geq N_{\varepsilon}$, $i\in S^{n}$, $\mu\in\mathcal{M}(A)$, we have
$\left|u_{i}^{n}(\mu,g^{n}_{-i})-u^{n}_{i}\Bigl{(}\mu,\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}\right|\leq\frac{\varepsilon}{4}.$
Similar to our proof of Lemma 2, we also use the sequence of random variables $\{x^{n}_{i}\}_{i\in I^{n},n\in\mathbb{Z}_{+}}$ to represent all the players’ strategies in $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. Let $x_{\mu}$ be a random variable that induces the distribution $\mu$, then we have
$u_{i}^{n}(\mu,g^{n}_{-i})=\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x_{\mu},\lambda^{n}(i)x_{\mu}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)x^{n}_{j}\Bigr{)}\Bigr{]},$
@@
$u^{n}_{i}\Bigl{(}\mu,\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}=\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\Bigr{)}\Bigr{]}.$
Hereafter, we restrict our attention to functions in $K_{\varepsilon}$. By equicontinuity, we know that for any $\varepsilon>0$, there exists $\eta>0$ such that for any $\tau,\widetilde{\tau}\in\mathcal{M}(A)$ with $\rho(\tau,\widetilde{\tau})\leq\eta$,
@@
Let $s(x_{\mu},x^{n}_{-i})(\omega)=\lambda^{n}(i)\delta_{x_{\mu}(\omega)}+\sum_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}(\omega)}$, and $s(\mu,g^{n}_{-i})=\lambda^{n}(i)\mu+\sum_{j\in I^{n}\backslash\{i\}}\\ \lambda^{n}(j)g^{n}(j)$, for all $\omega\in\Omega$, $\mu\in\mathcal{M}(A)$. The triangle inequality implies that,
@@
$+\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}$ (6)
By the definition of Prohorov metric $\rho$, we know that for any $\omega\in\Omega$, $\mu\in\mathcal{M}(A)$, $i\in I^{n}$,
$\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}\leq\sup\limits_{j\in I^{n}}\lambda^{n}(j).$
Since $\sup\limits_{j\in I^{n}}\lambda^{n}(j)\rightarrow 0$, there exists $N_{1}\in\mathbb{Z}_{+}$ such that for any $n\geq N_{1}$, we have $\mathop{\sup}\limits_{j\in I^{n}}\lambda^{n}(j)<\frac{\eta}{4}$. Hence for any $n\geq N_{1}$, $i\in I^{n}$, $\mu\in\mathcal{M}(A)$, $\omega\in\Omega$,
@@
By the same argument as above, we can see that for any $n\geq N_{1}$, $i\in I^{n}$, $\mu\in\mathcal{M}(A)$,
@@
Then we estimate $H_{1}^{(\frac{\eta}{2},n)}$ and $H_{2}^{(\frac{\eta}{2},n)}$ separately. Note that for any $n\geq N_{\varepsilon}$ and any player $i\in S^{n}$, we have $u_{i}^{n}\in K_{\varepsilon}$.
@@
$\left|\mathbb{E}\Bigl{[}u^{n}_{i}\bigl{(}x_{\mu},\lambda^{n}(i)\mu+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)g^{n}(j)\bigr{)}\Bigr{]}-\mathbb{E}\Bigl{[}u^{n}_{i}\bigl{(}x_{\mu},\lambda^{n}(i)\delta_{x_{\mu}}+\sum\limits_{j\in I^{n}\backslash\{i\}}\lambda^{n}(j)\delta_{x^{n}_{j}}\bigr{)}\Bigr{]}\right|\leq\frac{\varepsilon}{4}.$

### v2

@@ 4.3 Related literature
Substantial progress has been made in the theory of large games since the foundational works of Schmeidler (1973) and Mas-Colell (1984). More recent contributions include Kalai (2004), Sun (2006), Khan et al. (2013), Qiao and Yu (2014), Yu (2014), Deb and Kalai (2015), He et al. (2017), Khan et al. (2017), Kalai and Shmaya (2018), Khan et al. (2020), Sun et al. (2020), Carmona and Podczeck (2022), Chen et al. (2022), Hellwig (2022), Wu (2022), Yang (2022), and Chen et al. (2025). Significant advances has also been made in the analysis of large economies; see, for instance, Yannelis (2009), He and Yannelis (2016), and Anderson et al. (2022a, b). This section reviews literature most directly related to our analysis.
@@
This paper builds on prior work on symmetric equilibria in large games, notably Mas-Colell (1984), Khan and Sun (1995a, b), and Sun et al. (2020). Mas-Colell (1984) established the existence of pure strategy symmetric equilibria under the atomless condition and finite action sets. Khan and Sun (1995a, b) extended this result to games with countable action sets. More recently, Sun et al. (2020) considered randomized strategy symmetric equilibria and showed the existence of such equilibria in general large games. These results provide the theoretical foundation for our analysis. In contrast to most of this literature, which focuses on a continuum of players, we examine games with a large but finite population.
Our approach is also related to Kalai and Shmaya (2018), who studied incomplete-information repeated games with a large but finite number of players. To address analytical complexity, they introduced the notion of a (symmetric) imagined-continuum equilibrium, in which the societal summary—referred to as the “macro-strategy”—was unaffected by individual deviations. They established the existence of a symmetric Markov equilibrium in the imagined-continuum game and showed that it induced an approximate equilibrium in the finite-player game under a Lipschitz continuity condition on the outcome-generating function. While both papers examined symmetric equilibria in large finite games, key differences emerged. Kalai and Shmaya (2018) adopted symmetry as a simplifying assumption by considering homogeneous players with common payoffs and priors. This ensured that the same strategy applied in both continuum and finite-player settings, rendering the continuity condition in Theorem 1 trivially satisfied. In contrast, our model allows for heterogeneous players, rendering the continuity condition on the auxiliary mapping essential (see Example 1). Moreover, whereas Kalai and Shmaya (2018) assumed a finite action set and Lipschitz continuous payoffs, our framework accommodates general action spaces without these restrictions.
Theorem 2 relates to Kalai (2004), who studied the approximate ex post stability of Nash equilibria in Bayesian games with a large but finite number of players. He showed that if payoff functions satisfy certain equicontinuity conditions, then realizations of exact equilibria in the large finite game are approximately stable with high probability. The key distinction is that Theorem 2 considers a Nash equilibrium of the limiting large game $G$, which may not correspond to an exact Nash equilibrium in each finite-player game $G_{n}$. Moreover, as illustrated in Example 1, the continuity assumption on $\overline{g}$ in Theorem 2 is essential.
Finally, the paper contributes to the literature on the convergence of equilibria in large games. Earlier work (e.g., Green (1984), Housman (1988)) examined whether equilibrium correspondences are upper hemicontinuous as the number of players grows. More recent developments include Khan et al. (2013), Qiao and Yu (2014), and He et al. (2017). These papers focus primarily on pure strategies. We extend existing results by analyzing the convergence of randomized strategies, a technically distinct and less developed area in the study of large games.
@@ 5.1 Technical preparations
Let $A$ be a compact metric space with metric $d_{A}$, endowed with its Borel $\sigma$-algebra $\mathcal{B}(A)$. In this subsection, we introduce two equivalent metrics on the space of probability measures $\mathcal{M}(A)$. Using these metrics, we show that the distance between the sequence of societal summaries and the sequence of their realizations converges to zero in probability.
@@
It is known in the literature that $\rho$ and $\beta$ are equivalent metrics; see, for example, Bogachev (2007, Theorem 8.3.2). Recall that given a sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$, where each $g^{n}$ is a randomized strategy profile in the game $G_{n}$, there exists a probability space $(\Omega,\Sigma,\mathbb{P})$ and a sequence of random variables $\{x^{n}_{i}\}_{i\in I_{n},\,n\in\mathbb{Z}_{+}}$ mapping from $\Omega$ to $A$ such that:
@@
    For each $i\in I_{n}$ and $n\in\mathbb{Z}_{+}$, the distribution of $x^{n}_{i}$ equals $g^{n}(i)$;
@@
    For each $n\in\mathbb{Z}_{+}$, the random variables $\{x^{n}_{i}\}_{i\in I_{n}}$ are pairwise independent.
@@
Let $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite-player games, and let $g^{n}$ be a randomized strategy profile of $G_{n}$ for each $n\in\mathbb{Z}_{+}$. For each $\omega\in\Omega$, let $s(x^{n})(\omega)=\sum_{i\in I_{n}}\delta_{x^{n}_{i}(\omega)}\lambda_{n}(i)$ denote the realized societal summary associated with the strategy profile $g^{n}$. Thus $s(x^{n})$ defines a random variable from $(\Omega,\Sigma,\mathbb{P})$ to $\mathcal{M}(A)$. Then,
$\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0\quad\text{and}\quad\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0\quad\text{in probability,}$
where $s(g^{n})=\int_{I_{n}}g^{n}(i)\,\mathrm{d}\lambda_{n}(i)$.
@@
We divide the proof into two steps. In step 1, we show that for any bounded continuous function $h\colon A\to\mathbb{R}$ with $h||_{BL}\leq 1$, $\int_{A}h{\rm{d}}\bigl{(}{s(x^{n})-s(g^{n})}\bigr{)}\to 0$ in probability. In step 2, we show that $\beta\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability. Finally, by the equivalence of $\rho$ and $\beta$, we obtain that $\rho\bigl{(}s(x^{n}),s(g^{n})\bigr{)}\to 0$ in probability.
@@
$\int_{A}h{\rm{d}}\bigl{(}s(x^{n})-s(g^{n})\bigr{)}=\sum\limits_{i\in I_{n}}\lambda_{n}(i)\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}-\mathbb{E}\bigl{[}h\bigl{(}x^{n}_{i}\bigr{)}\bigr{]}\Bigr{)}\to 0$ (1)
in probability. Fix any $n\in\mathbb{Z}_{+}$, since $\{x^{n}_{i}\}_{i\in I_{n}}$ are pairwise independent and $h$ is a bounded continuous function, we know that $\{h(x^{n}_{i})\}_{i\in I_{n}}$ are also pairwise independent. By the definition of $\|h\|_{BL}\leq 1$, we have $\|h\|_{\infty}\leq 1$ and hence $1\leq h(x^{n}_{i})\leq 1$, $\text{var}\bigl{(}h(x^{n}_{i})\bigr{)}\leq 1$, for all $i\in I_{n}$. Moreover, by the independence of $\{h(x^{n}_{i})\}_{i\in I_{n}}$, we have
$\text{var}\Bigl{(}\sum\limits_{i\in I_{n}}h(x^{n}_{i})\lambda_{n}(i)\Bigr{)}=\sum\limits_{i\in I_{n}}\bigl{(}\lambda_{n}(i)\bigr{)}^{2}\text{var}\bigl{(}h(x^{n}_{i})\bigr{)}.$
Since $\mathbb{E}\bigl{[}\sum\limits_{i\in I_{n}}\lambda_{n}(i)h(x^{n}_{i})\bigr{]}=\sum\limits_{i\in I_{n}}\lambda_{n}(i)\mathbb{E}\bigl{[}h(x^{n}_{i})\bigr{]}$, for any $\varepsilon>0$, we have
$\mathbb{P}\Bigl{(}\Bigl{|}\sum_{i\in I_{n}}\lambda_{n}(i)\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}-\mathbb{E}\bigl{[}h\bigl{(}x^{n}_{i}\bigr{)}\big{]}\Bigr{)}\Bigr{|}\leq\varepsilon\Bigr{)}$
$\geqslant 1-\frac{\sum\limits_{i\in I_{n}}\bigl{(}\lambda_{n}(i)\bigr{)}^{2}\text{var}\Bigl{(}h\bigl{(}x^{n}_{i}\bigr{)}\Bigr{)}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\sum\limits_{i\in I_{n}}\bigl{(}\lambda_{n}(i)\bigr{)}^{2}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\mathop{\sup}\limits_{j\in I_{n}}\lambda_{n}(j)\sum\limits_{i\in I_{n}}{\lambda_{n}(i)}}{\varepsilon^{2}}$
$\geqslant 1-\frac{\mathop{\sup}\limits_{j\in I_{n}}\lambda_{n}(j)}{\varepsilon^{2}},$ (2)
where the first inequality is due to the Chebyshev’s inequality, and the last inequality follows from the fact that $\sum_{i\in I_{n}}{\lambda_{n}(i)}=1$. Combining with $\sup_{j\in I_{n}}\lambda_{n}(j)\to 0$ as $n\to\infty$, we can finish the proof of Formula (1).
@@
$\sum\limits_{i\in I_{n}}\lambda_{n}(i)\Bigl{(}p_{l}(x^{n}_{i})-\mathbb{E}\bigl{[}p_{l}(x^{n}_{i})\bigr{]}\Bigr{)}\rightarrow 0$ (3)
@@ 5.2 Proof of Lemma 1
Let $s(G,g)=\int_{I}\delta_{G(i)}\otimes g(i)\operatorname{d\!}\lambda(i)$ be the joint distribution of $G$ and $g$ on the product space $\mathcal{U}_{A}\times A$. Clearly, the marginal distribution $s(G,g)|_{\mathcal{U}_{A}}=\lambda G^{-1}$. Since both $\mathcal{U}_{A}$ and $A$ are Polish spaces, there exists a family of Borel probability measures $\{\mathcal{S}(u,\cdot)\}_{u\in\mathcal{U}_{A}}$ in $\mathcal{M}(A)$, which represents the disintegration of $s(G,g)$ with respect to $\lambda G^{-1}$ on $\mathcal{U}_{A}$. Let $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ be defined such that $\overline{g}(u)=\mathcal{S}(u)$ for all $u\in\mathcal{U}_{A}$. According to Sun et al. (2020, Lemma 5), the composition mapping $\widetilde{g}=\overline{g}\circ G$ is a Nash equilibrium of $G$ that satisfies $s(\widetilde{g})=s(g)=\tau$. Furthermore, we can ensure that for any player $i\in I$, her strategy $\widetilde{g}(i)$ is a best response with respect to the society summary s(g~)s($\widetilde${g}). This requirement can be met by modifying the strategies of a subset of players with measure zero.
@@ 5.3 Proof of Theorem 1
[…]
Step 1. We show that for any $\varepsilon>0$, there exist a sequence of subsets $S_{n}\subseteq I_{n}$ such that $\lambda_{n}(S_{n})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$, and $\{u_{i}^{n}\}_{i\in S_{n},n\in\mathbb{Z}_{+}}$ are equicontinuous and uniformly bounded by a constant $M_{\varepsilon}$. For simplicity, let $\mathcal{W}^{n}=\lambda_{n}{G_{n}}^{-1}$, and $\mathcal{W}=\lambda G^{-1}$. Since $A\times\mathcal{M}(A)$ is a compact metric space, the space of bounded and continuous functions $\mathcal{U}_{A}$ on $A\times\mathcal{M}(A)$ is a Polish space. By using the Prohorov theorem (Billingsley (1999, Theorem 5.2)), we know that $\{\mathcal{W}^{n}\}_{n\in\mathbb{Z}_{+}}$ is tight, which means that for any $\varepsilon>0$, there exists a compact set $K_{\varepsilon}\subset\mathcal{U}_{A}$ such that $\mathcal{W}^{n}(K_{\varepsilon})>1-\frac{\varepsilon}{2}$ for all $n\in\mathbb{Z}_{+}$.
Since $K_{\varepsilon}$ is a compact set that consists of bounded and continuous functions, the Arzelà-Ascoli theorem (Munkres (2000, Theorem 45.4)) implies that all the functions in $K_{\varepsilon}$ are uniformly bounded and equicontinuous functions. Let $M_{\varepsilon}$ denote a bound of all the functions in $K_{\varepsilon}$, and $S_{n}=\{i\in I_{n}|u_{i}^{n}\in K_{\varepsilon}\}$ for all $n\in\mathbb{Z}_{+}$. It is clear that $\lambda_{n}(S_{n})>1-\frac{\varepsilon}{2}$.
[…]
Similar to the proof of Lemma 2, we represent the strategies of all players in the sequence of games $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ using random variables $\{x^{n}_{i}\}_{i\in I_{n},n\in\mathbb{Z}_{+}}$, where each $\{x^{n}_{i}(\omega)\}_{i\in I_{n},n\in\mathbb{Z}_{+}}$ is a realization of a corresponding random variable $\{g_{i}^{n}\}_{i\in I_{n}}$. Let $x_{\mu}$ be a random variable that induces the distribution $\mu$, then we have
[…]
@@
$u^{n}_{i}\Bigl{(}\mu,\lambda_{n}(i)\mu+\sum\limits_{j\in I_{n}\backslash\{i\}}\lambda_{n}(j)g^{n}(j)\Bigr{)}=\mathbb{E}\Bigl{[}u_{i}^{n}\Bigl{(}x_{\mu},\lambda_{n}(i)\mu+\sum\limits_{j\in I_{n}\backslash\{i\}}\lambda_{n}(j)g^{n}(j)\Bigr{)}\Bigr{]}.$
Hereafter, we restrict our attention to functions $u,u_{i}^{n}\in K_{\varepsilon}$. By uniformly equicontinuity, we know that for any $\varepsilon>0$, there exists $\eta>0$ such that for any $\tau,\widetilde{\tau}\in\mathcal{M}(A)$ with $\rho(\tau,\widetilde{\tau})\leq\eta$,
@@
[…]
@@
$+\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}.$ (6)
By the definition of Prohorov metric $\rho$, we know that for any $\omega\in\Omega$, $\mu\in\mathcal{M}(A)$, $i\in I_{n}$,
$\rho\bigl{(}s(x^{n})(\omega),s(x_{\mu},x^{n}_{-i})(\omega)\bigr{)}\leq\sup\limits_{j\in I_{n}}\lambda_{n}(j).$
Since $\sup\limits_{j\in I_{n}}\lambda_{n}(j)\rightarrow 0$, there exists $N_{1}\in\mathbb{Z}_{+}$ such that for any $n\geq N_{1}$, we have $\mathop{\sup}\limits_{j\in I_{n}}\lambda_{n}(j)<\frac{\eta}{4}$. Hence for any $n\geq N_{1}$, $i\in I_{n}$, $\mu\in\mathcal{M}(A)$, $\omega\in\Omega$,
@@
By the same argument as above, we can see that for any $n\geq N_{1}$, $i\in I_{n}$, $\mu\in\mathcal{M}(A)$,
@@
Then we estimate $H_{1}^{(\frac{\eta}{2},n)}$ and $H_{2}^{(\frac{\eta}{2},n)}$ separately. Note that for any $n\geq N_{\varepsilon}$ and any player $i\in S_{n}$, we have $u_{i}^{n}\in K_{\varepsilon}$.
@@
[…]